# MindArm: Mechanized Intelligent Non-Invasive Neuro-Driven Prosthetic Arm System

### v1

@@ Abstract
Currently, people with disability or difficulty to move their arms (referred to as “patients”) have very limited technological solutions to efficiently address their physiological limitations. It is mainly due to two reasons: (1) the non-invasive solutions like mind-controlled prosthetic devices are typically very costly and require expensive maintenance; and (2) other solutions require costly invasive brain surgery, which is high risk to perform, expensive, and difficult to maintain. Therefore, current technological solutions are not accessible for all patients with different financial backgrounds. Toward this, we propose a low-cost technological solution called MindArm, a mechanized intelligent non-invasive neuro-driven prosthetic arm system. Our MindArm system employs a deep neural network (DNN) engine to translate brain signals into the intended prosthetic arm motion, thereby helping patients to perform many activities despite their physiological limitations. Here, our MindArm system utilizes widely accessible and low-cost surface electroencephalogram (EEG) electrodes coupled with an Open Brain Computer Interface and UDP networking for acquiring brain signals and transmitting them to the compute module for signal processing. In the compute module, we run a trained DNN model to interpret normalized micro-voltage of the brain signals, and then translate them into a prosthetic arm action via serial communication seamlessly. The experimental results on a fully working prototype demonstrate that, from the three defined actions, our MindArm system achieves positive success rates, i.e., 91% for idle/stationary, 85% for shake hand, and 84% for pick-up cup. This demonstrates that our MindArm provides a novel approach for an alternate low-cost mind-controlled prosthetic devices for all patients.
@@
EEG, Myoelectric, RPNI, BCI, Artefacts, Artificial Intelligence, Fast Fourier Transform.
@@ I Introduction
In recent years, approximately 5.4 million people live with paralysis in the United States alone [1] and 57.7 million people live with limb amputations globally [2]; see the global distribution of age-standardized amputation rates in Fig. 1. Among them, there are people with disability or difficulty to move their arms, which we refer to as “patients” in this paper for brevity. These patients have very limited technological solutions for efficiently addressing their physiological limitations as current technological solutions are not easily accessible for all patients with different financial backgrounds, because the existing solutions are typically very costly. For instance, the non-invasive solutions such as mind-controlled prosthetic devices are typically very costly (i.e., over \$100K), and hence expensive to maintain [3][4]. Other solutions require invasive brain surgery which is very expensive, high risk, and difficult to maintain, thereby hindering its widespread adoption [5][6]. Hence, there is a significant need for an alternative low-cost solution that help the patients to perform diverse activities.
@@ I-A State-of-the-Art Solutions and Their Limitations
To address the physiological limitations of the patients, the existing solutions can be loosely classified into two categories: (1) non-invasive solutions, such as mind-controlled prosthetic devices; (2) and invasive solutions, such as brain-computer interface (BCI).
@@ I-C Our Novel Contributions
Key Results: To evaluate our MindArm, we build a complete setup encompassing EEG acquisition module, compute module (i.e., DNN engine), and actuator module (e.g., servo motor and prosthetic arm), whose total cost is $\sim$\$450.
@@
Figure 3: MindArm Methodology: (a) Data Collection: EEG voltages are transmitted via Bluetooth to the GUI, which processes the data through a Fast Fourier Transform (FFT). The decomposed data are then saved to the cloud for further processing. (b) Neural Network Training: The dataset undergoes cleaning and is categorized into three groups. To accommodate UDP communication latency, the sampling rate is adjusted to 40 Hz. A window size of 80 samples, including a feature set of 20, is selected for training the network, which is subsequently saved to the cloud. (c) Deployment: The trained network interfaces with the prosthetic, designed with servo motors to enable 3 Degrees of Freedom (3DoF) movement. Network outputs are converted into servo actuations to perform the intended tasks.
@@ II-A EEG Data Collection & Extraction
The OpenBCI Ganglion board features 4 EEG channels and 2 references, facilitating the use of both dry comb electrodes and flat electrodes with a snap connection interface coated in silver-silver chloride [16]. Flat electrodes are positioned at Fp1 and Fp2 locations in the nasion, and comb electrodes at T3 and T4, with references placed at A1 and A2. The placement at Fp1 and Fp2 aims to enhance alpha wave detection. Although alpha waves are more prominent in the occipital lobe, placement near the inion (back of the head) often encounters greater noise. Consequently, while O1 and O2 locations yield stronger alpha and theta wave signals, achieving stable amplitudes across different frequencies using Fast Fourier Transform (FFT) analysis proves more challenging. Evidently, a scientific challenge arises when attempting to interpret data from only 4 EEG channels compared to numerous EEG channels in non-invasive state of the art devices.
@@
TABLE I: Comparison of EEG Devices
@@
Ganglion BCI \$400 4 channels None
@@
In the initial phase of the study, rather than directly decomposing the EEG signal into specific frequency bands, a more rudimentary threshold-based system was employed to ascertain the subject’s state of relaxation or concentration. The determination of the subject being relaxed or focused was then mapped to corresponding actions to be executed by the prosthetic device.
The results indicate promising trends; however, the current method presents a notable limitation: the states of concentration and relaxation are not distinctly categorized by Graphical User Interface (GUI). Consequently, the algorithm fails to differentiate between ‘relaxed’ and ‘concentrated’ states, only recognizing ‘relaxed and stationary’ as well as ‘concentrated and stationary’ states. This issue stems from an observable overlap in the metric thresholds for relaxation and concentration within the GUI. Therefore, it is imperative to develop an alternative approach that enables accurate prediction across all three desired states: handshaking, cup grasping, and remaining stationary. Notably, a parallel in amplitude across various frequencies is observed when comparing the thought of handshaking with the relaxed state, and the thought of cup grasping with the concentrated state.
@@
To enhance the accuracy of state prediction, the proposed methodology involves adopting a strategy of associating relaxation with the intent to shake hands and concentration with the intent to pick up a cup.
To validate the efficacy of this approach, a novel signal processing technique is employed, which decomposes EEG signals into five distinct brain wave categories: delta, theta, alpha, beta, and gamma, each within their characteristic frequency bands: delta (1–4 Hz), theta (4–8 Hz), alpha (8–13 Hz, with sub-bands alpha-1 8–10 Hz and alpha-2 11–13 Hz), beta (13–30 Hz), and gamma (above 30 Hz). This spectral decomposition allows for more precise extraction of features from each channel, thereby reducing noise transmission through the User Datagram Protocol (UDP) network and enhancing algorithmic accuracy. The choice of UDP is driven by its expedited data transfer capabilities, absence of connection establishment procedures, and greater efficiency via lower bandwidth usage and overhead, thus compensating for potential latency and ensuring a robust set of features for the neural network classifier, contributing to a low-maintenance system [17].
@@
Figure 11: Comparison of TCP and UDP Protocols. TCP initiates communication with a three-way handshake, starting with a synchronize (SYN) message from the client (a), followed by a synchronize-acknowledgment (SYN-ACK) from the server, and concluding with an acknowledgment (ACK) from the client (b), thereby establishing a connection. In contrast, UDP operates without a handshake, allowing data to be sent and received without establishing a persistent connection.
UDP Networking and Data Transmission # Python code for UDP networking UDP_IP = ”127.0.0.1” UDP_PORT = 4005 sock = socket.socket(socket.AF_INET, socket.SOCK_DGRAM) sock.bind((UDP_IP, UDP_PORT)) # Data is sent as a JSON message of the form: [ [delta_ch_1, theta_ch_1, alpha_ch_1, beta_ch_1, gamma_ch_1], [delta_ch_2, theta_ch_2, alpha_ch_2, beta_ch_2, gamma_ch_2], [delta_ch_3, theta_ch_3, alpha_ch_3, beta_ch_3, gamma_ch_3], [delta_ch_4, theta_ch_4, alpha_ch_4, beta_ch_4, gamma_ch_4], ] # Example received message with the IDE: received message: b’{”type”:”bandPower”,”data”:[Array[5x4]]}’
@@ II-B Neural Network Training
In the exploration of optimal neural network architectures for our dataset, a diverse range of models was assessed. These included simple Feedforward Neural Networks (FFNNs), Recurrent Neural Networks (RNNs), Long Short-Term Memory networks (LSTMs), CNN-LSTM hybrids, and Transformer-based networks. Among these, the Transformer-based network outperformed other models achieving an impressive validation accuracy of 97.1%. This superior performance can be attributed to the Transformer’s ability to process sequential data in parallel and its efficient handling of long-range dependencies, which are critical for understanding complex patterns within the EEG dataset. However, the trade-off for this high level of accuracy involves increased computational resources and training time, compared to simpler models like FFNNs or RNNs. The decision to employ a Transformer-based model thus reflects a strategic balance between seeking optimal performance and managing the computational costs associated with more sophisticated architectures. The transformer architecture is illustrated in Fig. 14. The training accuracy of these networks are depicted in Fig. 15 showcasing the comparative training performance across epochs for the various network architectures.
Figure 13: EEGTransformer Architecture. The depicted Transformer-based model processes segmented EEG data sequences, leveraging a multi-head self-attention mechanism and subsequent layer normalizations within the Transformer Encoder. Data flows through an average pooling layer before passing through two feed-forward networks, culminating in the classification of the prosthetic arm’s actions. These predicted actions are then communicated to the prosthetic arm via serial communication for execution.
@@
Accordingly, an optimized window size of 100 was implemented. This setup accumulates 100 rows of CSV data—each row representing a 20-dimensional vector from the EEG—into a single tensor. This tensor is then reshaped into a 1x2000 matrix (20x100), serving as the input for the neural network. This approach ensures that the input data effectively represents approximately two seconds of EEG data, allowing for more accurate and representative model outputs as live EEG data is streamed.
@@
TABLE II: Dataset Dimensions Before and After Processing
@@
With the refactored algorithm, an output of the intended user action is transmitted to the Arduino during inference every 2 seconds. This decreases volatility of the actions as a substantial greater number of features is taking into consideration for a sustained time period, thereby enhancing the robustness of the model.
Additionally, the neural network undergoes pre-training prior to its deployment and is subsequently stored on a cloud service. This strategy ensures that the duration for outputting a label is shorter than the interval between consecutive inputs, thereby reducing the potential for data loss during the wireless EEG streaming process.
@@ II-C Prosthetic Design and Label Feedback
The prosthetic control system transmits the output label to the Arduino via serial communication every 2 seconds. At a standardized baud rate, the Arduino interprets the received number ‘0, 1, or 2’ and initiates the corresponding action. To prevent potential damage to the servos and artificial tendons due to rapid oscillation between positions, the algorithm is designed to pause reading incoming data until the current action is fully executed.
@@
Initially, the prosthetic’s actuation mechanism relied on the contraction and relaxation of tendons, facilitated by a servo horn. However, this design was found impractical due to the persistent friction between the tendons, made of braided fishing line, and the joint pins. Consequently, this friction led to the tendons’ degradation over time, resulting in both wear and tear of the prosthetic components and a decline in performance.
@@ III-A Experimental Setup
The experimental setup for the prosthetic system comprises an array of specialized tools and technologies. The neural signal acquisition is managed using the Ganglion board and accompanying GUI from OpenBCI. Data communication is facilitated via UDP networking. For the development and training of the neural network, the PyTorch framework is employed. The system’s commands are transmitted through serial communication with the Arduino C++ IDE. The design and prototyping phase utilizes Tinkercad and Fusion360 for 3D modeling, alongside custom settings in Prusa Slicer for 3D printing and circuit schematic refinement.
@@ III-B Analysis
TABLE III: Classification Report
@@ IV Conclusion
In this paper, we present a MindArm methodology to realize a low-cost mind-controlled prosthetic arm solution for people of determination, so that they can move their hands for doing activities. It translates brain signals into the intended arm motion by utilizing EEG technologies and employing DNN model to interpret brain signals into a prosthetic arm action. The experimental results show that, our MindArm system achieves positive success rates in three different actions, i.e., 90% for idle/stationary, 80% for shake hand, and 80% for pick-up cup. This demonstrates that our MindArm provides a novel approach for an alternate low-cost mind-controlled prosthetic devices for all people.

### v2

@@ Abstract
Currently, individuals with arm mobility impairments (referred to as “patients”) face limited technological solutions due to two key challenges: (1) non-invasive prosthetic devices are often prohibitively expensive and costly to maintain, and (2) invasive solutions require high-risk, costly brain surgery, which can pose a health risk. Therefore, current technological solutions are not accessible for all patients with different financial backgrounds. Toward this, we propose a low-cost technological solution called MindArm, an affordable, non-invasive neuro-driven prosthetic arm system. MindArm employs a deep neural network (DNN) to translate brain signals, captured by low-cost surface electroencephalogram (EEG) electrodes, into prosthetic arm movements. Utilizing an Open Brain Computer Interface and UDP networking for signal processing, the system seamlessly controls arm motion. In the compute module, we run a trained DNN model to interpret filtered micro-voltage brain signals, and then translate them into a prosthetic arm action via serial communication seamlessly. Experimental results from a fully functional prototype show high accuracy across three actions, with 91% for idle/stationary, 85% for handshake, and 84% for cup pickup. The system costs approximately \$500-550, including \$400 for the EEG headset and \$100-150 for motors, 3D printing, and assembly, offering an affordable alternative for mind-controlled prosthetic devices.
@@
EEG, OpenBCI, non-invasive, prosthetic arm, deep learning, low cost.
@@ I Introduction
Globally, millions of people live with disabilities, including limb amputations, which severely limit their ability to perform everyday tasks. In recent years, approximately 5.4 million people live with paralysis in the United States alone [1] and 57.7 million people live with limb amputations globally [2]; see the global distribution of age-standardized amputation rates in Fig. 1. Among them, there are people with disability or difficulty to move their arms, which we refer to as “patients” in this paper for brevity. These patients have limited access to effective technological solutions for managing their physiological limitations, primarily because many current innovations are not affordable for individuals from diverse financial backgrounds.
Existing prosthetic solutions, particularly mind-controlled devices, are prohibitively expensive, often exceeding \$100k [3, 4]. Invasive solutions also exist but carry additional costs and risks associated with surgery [5, 6]. To address the need for a cost-effective, non-invasive solution, we propose MindArm, a system that utilizes widely available surface EEG electrodes to control a prosthetic arm via a DNN engine.
Hence, there is a significant need for an alternative low-cost solution that help the patients to perform diverse activities.
@@ I-A State-of-the-Art Solutions and Their Limitations
Various state-of-the-art prosthetic solutions exist, but they are typically expensive and not accessible to all patients. Below is a comparison between MindArm and leading commercial prosthetic arms.
TABLE I: Comparison of MindArm with Other Prosthetic Solutions
Prosthetic Arm Cost (USD) Features/ Maintenance
MindArm (ours) 500 - 550 EEG-based, 3D-printed, Low
DEKA Arm System 100K - 200K EMG-based, High
Ottobock 60K - 120K Advanced bionic arm, Medium
Bebionic Arm 25K - 75K Multi-articulating hand, Medium
i-Limb Quantum 40K - 60K Advanced multi-grip, Medium
To address the physiological limitations of the patients, the existing solutions can be loosely classified into two categories: (1) non-invasive solutions, such as surface mounted eeg electrode devices; (2) and invasive solutions, such as implantable brain-computer interface (BCI).
@@ I-C Our Novel Contributions
Figure 3: MindArm Methodology: (a) Data Collection: EEG voltages are transmitted via Bluetooth to the GUI, which processes the data through a Fast Fourier Transform (FFT). The decomposed data are then saved to the cloud for further processing. (b) Neural Network Training: The dataset undergoes cleaning and is categorized into three groups. To accommodate UDP communication latency, the sampling rate is adjusted to 40 Hz. A window size of 80 samples, including a feature set of 20, is selected for training the network, which is subsequently saved to the cloud. (c) Deployment: The trained network interfaces with the prosthetic, designed with servo motors to enable 4 Degrees of Freedom (4DoF) movement. Network outputs are converted into servo actuations to perform the intended tasks.
Key Results: To evaluate our MindArm, we build a complete setup encompassing EEG acquisition module, compute module (i.e., DNN engine), and actuator module (e.g., servo motor and prosthetic arm) allowing for for 3 degrees of freedom, whose total cost is $\sim$\$550.
@@ II-A EEG Data Collection & Extraction
The OpenBCI Ganglion board features 4 EEG channels and 2 references, facilitating the use of both dry comb electrodes and flat electrodes with a snap connection interface coated in silver-silver chloride [16]. Flat electrodes are positioned at Fp1 and Fp2 locations in the nasion, and comb electrodes at T3 and T4, with references placed at A1 and A2. The placement at Fp1 and Fp2 aims to enhance alpha wave detection. Although alpha waves are more prominent in the occipital lobe, placement near the inion (back of the head) often encounters greater noise. Consequently, while O1 and O2 locations yield stronger alpha and theta wave signals, achieving stable amplitudes across different frequencies using Fast Fourier Transform (FFT) analysis proves more challenging. Evidently, a scientific challenge arises when attempting to interpret data from only 4 EEG channels compared to numerous EEG channels in non-invasive state of the art devices. Fig. 4 & 5 shows user wearing the headset.
@@
Figure 5: 10-20 system for electrode placement, highlighting the region FP1, FP2, T3, and T4 being used for the measurements.
TABLE II: Comparison of EEG Devices
@@
Ganglion BCI (ours) \$400 4 channels None
@@
Figure 7: The left panel displays EEG band power across Delta, Theta, Alpha, Beta, and Gamma frequencies. The right panel (Focus Widget) presents the subject’s cognitive state. The large circle indicates the overall focus status, where ”not focused” is shown. Adjacent bar graphs depict the threshold comparison for Alpha and Beta waves, indicating the subject’s relaxation (Alpha) and concentration (Beta) levels relative to preset thresholds (yellow line), helping classify the subject’s mental state.
In the initial phase of the study, rather than directly decomposing the EEG signal into specific frequency bands, as shown in Fig. 6, a more rudimentary threshold-based system [17] was employed to ascertain the subject’s state of relaxation or concentration, as previously used in EEG signal processing for cognitive state detection. The determination of the subject being relaxed or focused was then mapped to corresponding actions to be executed by the prosthetic.
The results indicate promising trends; however, the current method presents a notable limitation: the states of concentration and relaxation are not distinctly categorized by Graphical User Interface (GUI) shown in Fig. 7. Consequently, the algorithm fails to differentiate between ‘relaxed’ and ‘concentrated’ states, only recognizing ‘relaxed and stationary’ as well as ‘concentrated and stationary’ states. This issue stems from an observable overlap in the metric thresholds for relaxation and concentration within the GUI. Therefore, it is imperative to develop an alternative approach that enables accurate prediction across all three desired states: handshaking, cup grasping, and remaining stationary. Notably, a parallel in amplitude across various frequencies is observed when comparing the thought of handshaking with the relaxed state, and the thought of cup grasping with the concentrated state as shown in Fig. 8.
@@
To enhance the accuracy of state prediction, the proposed methodology involves adopting a strategy of associating relaxation with the intent to shake hands and concentration with the intent to pick up a cup. See Fig. 10 & 11.
To validate the efficacy of this approach, a novel signal processing technique is employed, which decomposes EEG signals into five distinct brain wave categories: delta, theta, alpha, beta, and gamma, each within their characteristic frequency bands: delta (1–4 Hz), theta (4–8 Hz), alpha (8–13 Hz, with sub-bands alpha-1 8–10 Hz and alpha-2 11–13 Hz), beta (13–30 Hz), and gamma (above 30 Hz) as illustrated in Fig. 9. This spectral decomposition allows for more precise extraction of features from each channel, thereby reducing noise transmission through the User Datagram Protocol (UDP) network and enhancing algorithmic accuracy. The choice of UDP is driven by its expedited data transfer capabilities, absence of connection establishment procedures, and greater efficiency via lower bandwidth usage and overhead, thus compensating for potential latency and ensuring a robust set of features for the neural network classifier, contributing to a low-maintenance system [18].
@@ II-B Neural Network Training
In the exploration of optimal neural network architectures for our dataset, a diverse range of models was assessed. These included simple Feedforward Neural Networks (FFNNs), Recurrent Neural Networks (RNNs), Long Short-Term Memory networks (LSTMs), CNN-LSTM hybrids, and Transformer-based networks. Among these, the Transformer-based network outperformed other models achieving an impressive validation accuracy of 97.1%. This superior performance can be attributed to the Transformer’s ability to process sequential data in parallel and its efficient handling of long-range dependencies, which are critical for understanding complex patterns within the EEG dataset. However, the trade-off for this high level of accuracy involves increased computational resources and training time, compared to simpler models like FFNNs or RNNs. The decision to employ a Transformer-based model thus reflects a strategic balance between seeking optimal performance and managing the computational costs associated with more sophisticated architectures. The transformer architecture is illustrated in Fig. 14. The training accuracy of these networks are depicted in Fig. 15 showcasing the comparative training performance across epochs for the various network architectures. The transformer model showcases the highest validation accuracy. Fig. 16 shows window size optimization.
Figure 14: EEG Transformer Architecture. The depicted Transformer-based model processes segmented EEG data sequences, leveraging a multi-head self-attention mechanism and subsequent layer normalization within the Transformer Encoder. Data flows through an average pooling layer before passing through two feed-forward networks, culminating in the classification of the prosthetic arm’s actions. These predicted actions are then communicated to the prosthetic arm via serial communication for execution.
@@
Accordingly, an optimized window size of 100 was implemented. This setup accumulates 100 rows of CSV data—each row representing a 20-dimensional vector from the EEG—into a single tensor. This tensor is then reshaped into a 1x2000 matrix (20x100), serving as the input for the neural network. This approach ensures that the input data effectively represents approximately two seconds of EEG data, allowing for more accurate and representative model outputs as live EEG data is streamed as shown in Fig. 17.
@@
TABLE III: Dataset Dimensions Before and After Processing
@@
The refactored algorithm transmits the user’s intended action to the Arduino every 2 seconds, improving model robustness by considering more features over a sustained period. Pre-training the neural network and storing it on the cloud ensures faster label output than the input interval, minimizing data loss during wireless EEG streaming.
@@ II-C Prosthetic Design and Label Feedback
The prosthetic control system transmits the output label to the Arduino via serial communication every 2 seconds. At a standardized baud rate, the Arduino interprets the received number ‘0, 1, or 2’ and initiates the corresponding action. To prevent potential damage to the servos and artificial tendons due to rapid oscillation between positions, the algorithm is designed to pause reading incoming data until the current action is at least a third way through executing.
@@
Initially, the prosthetic’s actuation mechanism relied on the contraction and relaxation of tendons, facilitated by a servo horn. However, this design was found impractical due to the persistent friction between the tendons, made of braided fishing line, and the joint pins. Consequently, this friction led to the tendons’ degradation over time, resulting in both wear and tear of the prosthetic components and a decline in performance. See Fig. 18 - 21.
@@ III-A Experimental Setup
The experimental setup includes the Ganglion board with OpenBCI GUI for neural signal acquisition and UDP networking for data communication. Neural network development is done using PyTorch, with commands transmitted via Arduino C++ IDE. Design and prototyping use Tinkercad, Fusion360 for 3D modeling, and Prusa Slicer for 3D printing and circuit refinement.
@@ III-B Analysis
TABLE IV: Classification Report
@@ IV Conclusion
In this paper, we introduced the MindArm methodology, a low-cost, non-invasive, mind-controlled prosthetic arm designed to assist individuals with disabilities in performing everyday tasks. The system translates brain signals into arm motion using EEG technology, with a deep neural network (DNN) interpreting these signals to control prosthetic actions. Experimental test results demonstrate high success rates: 90% for idle/stationary, 80% for shaking hands, and 80% for picking up a cup, validating the effectiveness of the system as an affordable alternative to existing prosthetic solutions shown in Fig. 22. While MindArm is cost-effective and accessible, it faces limitations such as noise in surface EEG signals and a restricted range of actions. The hardware, though affordable, may also have durability issues. Future efforts will focus on refining signal accuracy, expanding action capabilities, and enhancing hardware robustness through improved control mechanisms and signal filtering.
Acknowledgment
This work was partially supported by the NYUAD Center for Artificial Intelligence and Robotics (CAIR), funded by Tamkeen under the NYUAD Research Institute Award CG010.